OpenStudy (anonymous):

f(5)=0 and f(-4)=0 factor completely x^4-27x^2-14x+120

OpenStudy (anonymous):

it must factor as \[(x-5)(x+4)(something)\] i guess you can find the something either by division, or by thinking

OpenStudy (anonymous):

i can't write division here, but we can use the "think" method if you like

OpenStudy (anonymous):

\[x^4-27x^2-14x+120=(x-5)(x+4)(whatever)\] or multiplying out you get \[(x^2-x-20)(whatever)=x^4-27x^2-14x+120\]

OpenStudy (anonymous):

we know it has to be \[(x^2-x-20)(ax^2+bx+c)=x^4-27x^2-14x+120\] what we need are \(a,b,c\) \(a\) and \(c\) should be pretty obvious, especially \(a\) since you need \(x^4\) so \(a=1\)

OpenStudy (anonymous):

also the constant is \(120\) and you have a constant of \(-20\) so \(c=120\div(-20)=-6\)

OpenStudy (anonymous):

\[(x^2-x-20)(x^2+bx-6)=x^4-27x^2-14x+120\]

OpenStudy (anonymous):

the only real work is finding \(b\)

OpenStudy (loser66):